# EEGDM: Learning EEG Representation with Latent Diffusion Model

### v2

@@ Abstract
Index Terms:
Self-supervised learning, diffusion model, EEG representation learning.
I Introduction
@@
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies [mohsenvand2020contrastive]. Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components [zander2011towards].
@@
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data [zhao2021plug, zheng2023scatterformer, chen2025long]. Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, three large EEG models, EEGPT [wangeegpt], LaBraM [jiang2024large], and CBraMod [wang2024cbramod] have recently been proposed. These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
@@
In this paper, we introduce the Diffusion Model (DM) [ho2020denoising, dhariwal2021diffusion] into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution [ho2020denoising]. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
@@
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity [gramfort2013time, cole2019cycle]. To address these challenges, we incorporate two key technical components in our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones [sanei2013eeg]. This transformation mitigates the challenge of reconstructing the original noisy signal. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) [rombach2022high] and Diffusion Transformers (DiTs) [peebles2023scalable]) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks and exhibiting strong generalizability.
@@
The remainder of this work is organized as follows. First, we describe the related works in Section II. Next, the proposed algorithm is introduced in detail in Section III. After that, experimental results are shown in Section IV. Finally, the conclusion is given in Section V.
II Related work
II-A Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision [kenton2019bert, he2022masked]. However, its full potential in EEG analysis remains largely unexplored [jiang2024neurolm].
Some studies [mohsenvand2020contrastive, banville2021uncovering] have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework [chen2020simple] to time series data. Another line of work extends the mask idea to EEG analysis [mohammadi2024eeg2rep]. For instance, Chien et al. [chien2022maeeg] propose a masked auto-encoder, which learns the EEG representation by learning to reconstruct the masked EEG features using a Transformer architecture. BENDR [kostas2021bendr] uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM [jiang2024large] pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT [wangeegpt] introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features. CBraMod [wang2024cbramod] introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
@@ II-B Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. [sohl2015deep] and subsequent advances by Ho et al. [ho2020denoising] established diffusion models as powerful tools, leading to widespread applications in generative tasks [rombach2022high, peebles2023scalable, zhang2023adding]. More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features [mittal2023diffusion, chen2024deconstructing, hudson2024soda].
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation [huang2024eegdfus, chen2024improving, huang2025sad], or a few specific supervised tasks [kim2023diff, kim2024brain]. However, the potential of diffusion models for generic EEG representation learning remains largely unexplored. This presents a significant opportunity, as learning the complex temporal and spatial characteristics of EEG signals can benefit from the understanding of rich data distributions inherent in diffusion models’ denoising process. In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
@@
Figure 1: The overall architecture of EEGDM. The original EEG data is projected onto a latent space using PCA. Noise is added in this latent space, and the denoised latent representation is predicted by learning the conditional latent diffusion model. The prediction is then projected back to the original space by inverse PCA to obtain the generated target EEG signal. The encoder transforms the source EEG signal (including original data and augmentations) into a concise EEG representation. It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
III Method
@@
Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying properties, structures, and dynamics of the data during the generative process. Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
III-A Framework Architecture
@@
Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. It is worth noting that in the denoising process of DM, directly reconstructing raw EEG signals is challenging due to their inherently low SNR. To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR. The overall architecture of EEGDM is illustrated in Fig. 1.
@@
Figure 2: The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
III-B The EEG Encoder
@@
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do [hinton2006reducing, higgins2017beta], our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
@@
Our EEG encoder is based on the Vision Transformer (ViT) architecture [dosovitskiy2020image] that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work [mohsenvand2020contrastive] to be the most effective for extracting useful features.
@@
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation. As shown in Fig. 2, we train a linear predictor as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ task-specific loss functions for different downstream objectives. The implementation details can be found in the subsection IV-B.
III-C EEG Data Generation by Diffusion Model
@@
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we first enhance the SNR of the raw EEG signals using PCA. Similar to the LDMs [rombach2022high], we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
III-C1 Diffusion Formulation
@@
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (1)
@@
$p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$ (2)
@@
$\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$ (3)
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. [nichol2021improved]: train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
$\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$. (4)
III-C2 Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance [ho2022classifier] can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
$\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$ (5)

### v3

@@ Abstract.
Electroencephalography, Representation Learning, Self-supervised Learning, Diffusion Transformers
conference: ; ; copyright: noneisbn: 978-1-4503-XXXX-X/2018/06submissionid: 5469ccs: Computing methodologies Machine learningccs: Human-centered computing Human computer interaction (HCI)$c$s: Information systems Multimedia information systems
1. Introduction
@@
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al., 2020). Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe, 2011).
@@
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Jia et al., 2020, 2021; Jiang et al., 2024a). Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, three large EEG models, EEGPT (Wang et al., 2024), LaBraM (Jiang et al., 2024c), and CBraMod (Wang et al., 2025) have recently been proposed. These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
@@
In this paper, we introduce the Diffusion Model (DM) (Ho et al., 2020; Dhariwal and Nichol, 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction paradigms that primarily focus on local patches recovery, DM learns effective representations through a self-supervised denoising objective. This objective necessitates a profound understanding of the underlying data distribution, thereby compelling the latent space to preserve macro-level neural patterns. To successfully denoise a latent representation into a coherent EEG segment, the model must capture the holistic phase relationships and cross-channel correlations that define specific brain states. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
@@
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al., 2013; Cole and Voytek, 2019). To address these challenges, we incorporate two synergistic components into our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers, 2013). This transformation mitigates the challenge of reconstructing the original noisy signal. More importantly, PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views. Meanwhile, the diverse augmented inputs prevent the model from losing critical information when training against a reduced-dimension (PCA) target. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al., 2022) and Diffusion Transformers (DiTs) (Peebles and Xie, 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks and exhibiting strong generalizability.
@@
2. Related Work
2.1. Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova, 2019; He et al., 2022). However, its full potential in EEG analysis remains largely unexplored (Jiang et al., 2025).
Some studies (Mohsenvand et al., 2020; Banville et al., 2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al., 2020) to time series data. Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al., 2024). For instance, BENDR (Kostas et al., 2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM (Jiang et al., 2024c) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT (Wang et al., 2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features. CBraMod (Wang et al., 2025) introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
@@ 2.2. Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) and subsequent advances by Ho et al. (Ho et al., 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al., 2022; Peebles and Xie, 2023; Zhang et al., 2023). More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al., 2023; Chen et al., 2024b; Hudson et al., 2024).
@@
Figure 1. The overall architecture of EEGDM. The original EEG data is projected onto a latent space using PCA. Noise is added in this latent space, and the denoised latent representation is predicted by learning the conditional latent diffusion model. The prediction is then projected back to the original space by inverse PCA to obtain the generated target EEG signal. The encoder transforms the source EEG signal (including original data and augmentations) into a concise EEG representation. It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al., 2024; Chen et al., 2024a; Huang et al., 2025), or on a few specific supervised tasks (Kim et al., 2023, 2024). Recently, Puah et al. (Puah et al., 2025) proposed an EEG representation learning framework, also termed EEGDM. To avoid ambiguity, we refer to their method as Puah-EEGDM-SSM. Their approach integrates a Structured State-Space Model (SSM) into the diffusion pre-training paradigm to better capture the temporal dynamics of EEG signals. In contrast, our EEGDM adopts a fundamentally different architectural paradigm. Specifically, we employ a Latent Diffusion Transformer (DiT) operating on a refined latent manifold, which is specifically designed to capture global inter-channel spatial synchrony and complex phase relationships—critical neural features that are overlooked by the linear recurrence mechanisms inherent in SSM.
3. Method
@@
Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying distributions, structures, and dynamics of the data during the generative process. Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
3.1. Framework Architecture
@@
Our proposed self-supervised EEG representation learning framework, named EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. Notably, reconstructing raw EEG signals directly remains formidable due to their inherently low SNR. To mitigate this, we employ PCA to project raw data into a refined latent space with enhanced SNR, shifting the reconstruction target to these denoised components. By focusing on the top-K principal components, the diffusion module effectively constrains the search space to the most energetically significant neural manifolds, thereby alleviating the optimization burden. Crucially, while EEG augmentation introduces view diversity to force the extraction of invariant features, PCA provides a stable, denoised reconstruction objective. The overall architecture of EEGDM is shown in Fig. 1.
@@
3.2. The EEG Encoder
@@
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov, 2006; Higgins et al., 2017), our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
@@
Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy, 2021) that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al., 2020) to be the most effective for extracting useful features.
@@
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. As illustrated in Fig. 2, we evaluate the model’s performance on various downstream tasks by appending a task-specific linear prediction head to the pre-trained encoder. For each downstream objective, tailored loss functions are employed to ensure optimal task-specific adaptation. The implementation details are provided in Section 4.2.
Figure 2. The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
3.3. EEG Data Generation by Diffusion Model
@@
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we utilize PCA to project raw data into a high-SNR subspace, establishing a refined latent manifold as the denoising objective. Similar to the LDMs (Rombach et al., 2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
3.3.1. Diffusion Formulation
@@
(1) $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$
@@
(2) $p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$
@@
(3) $\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. (Nichol and Dhariwal, 2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
(4) $\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$.
3.3.2. Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance (Ho and Salimans, 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
(5) $\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$